OpenStudy (anonymous):

I need help on differentiate. I know to use quotient rule, but I can't get the correct answer? -2t^5-4t^3+10/(t^4-3t^3+2)^2 is what i got? y = t^2 + 1 divided by t^4 − 3t^2 + 2

OpenStudy (anonymous):

numerator was my attempt to simplify, i originally got 2t(t^4-3t^2+2)-(t^2+1)(4t^3-6t)

zepdrix (zepdrix):

Hmm your numerator looks good. I don't think you should try to simplify this one +_+ Ew..

zepdrix (zepdrix):

Oh I guess you get some nice cancellations n stuff if you do.. grr

OpenStudy (anonymous):

okay thank you zepdrix ill just let it go :)

OpenStudy (anonymous):

\[y=\frac{t^2+1}{t^4-3t^2+2}\] So \[\begin{align*}y'&=\frac{\left(t^4-3t^2+2\right)(2t)-\left(t^2+1\right)\left(4t^3-6t\right)}{\left(t^4-3t^2+2\right)^2}\\ &=\frac{2t\left[\left(t^4-3t^2+2\right)-\left(t^2+1\right)\left(2t^2-3\right)\right]}{\left(t^4-3t^2+2\right)^2}\\ &=\frac{2t\left[\left(t^4-3t^2+2\right)-\left(2t^4-t^2-3\right)\right]}{\left(t^4-3t^2+2\right)^2}\\ &=\frac{2t\left[-t^4-2t^2+5\right]}{\left(t^4-3t^2+2\right)^2}\\ &=\frac{-2t\left[t^4+2t^2-5\right]}{\left(t^4-3t^2+2\right)^2}\end{align*}\] is what I'm getting...